OpenStudy (math_bender_15):

The diameter of a construction pipe is 5ft . If the pipe is 20ft long, what is its volume?

OpenStudy (lgbasallote):

\[\large V = \pi (\frac{d}{2})^2 h\]

OpenStudy (anonymous):

v = pi r^2 h....r = 1/2 diameter and h is equal to your length

OpenStudy (lgbasallote):

you have diameter and height given so just plug in

OpenStudy (math_bender_15):

okay